A normally distributed population has mean 57.7 and standard deviation 12.1.

a. Find the probability that a single randomly selected element X of the population is less than 45

b. Find the mean and standard deviation of 𝑥̅for samples of size 16.

c. Find the probability that the mean of a sample of size 16 drawn from this population is less than 45.


1
Expert's answer
2021-05-07T08:42:49-0400

"\mu=57.7,\sigma=12.1"


(a) "P(X<45)=P(z<\dfrac{45-57.7}{12.1})=P(Z<-1.0409)=0.14695"


(b) Sample mean "\mu_X=\mu=57.7" , standard deviation "s=\dfrac{\sigma}{\sqrt{n}}=\dfrac{12.1}{\sqrt{16}}=\dfrac{12.1}{4}=3.025"


(c) "P(\bar{X}<45)=P(z<\dfrac{45-57.7}{3.025})=P(z<-4.19835)= 0.00001"